Eliminate the fractions in the following equation: x(3/4) + y(2/3) + z(5/6)= 2.
Answer by Alan3354(69443)

Eliminate the fractions in the following equation: x(3/4) + y(2/3) + z(5/6)= 2
Multiply the denominators
4*3*6 = 72
Multiply thru by 72
54x + 48y + 60z = 144
--------------------------
72 is not the LCD, but multiplying the DENs will always work, and takes less time to figure out.